Solve the system by the substitution method.
2x+y-x=3y+5
y+2=x-5

Answer by waynest(281)

use y + 2 = x - 5
y = x -2 -5
y = x - 7
2x + (x -7) -x = 3(x - 7) + 5
2x + x - 7 - x = 3x -21 + 5
2x - 7 = 3x - 16
16 -7 = 3x - 2x
9 = x
y = 9 - 7
y = 2
check:
2(9) + 2 - 9 = 3(2) + 5
18 + 2 - 9 = 6 + 5
20 - 9 = 11
11 = 11
2 + 2 = 9 - 5
4 = 4